Main INDEX
Monthly INDEX
PREV
NEXT
User name Raue
Log entry time 07:46:01 on June22,2005
Entry number 146907
keyword=shift summary
SHIFT SUMMARY
0:50 Started shift with problems of septum IOCs being down and replay not working. Might have figured out IOC issue. RC called in to fix replay problem.
1:27 Septa back and beam ready. Starting runs. Replay still a no go.
1:41 Beam has been pretty unstable for the last 15 minutes. Last run 4798 is probably crap.
1:45 Run 4799 started. All's well.
2:32 MCC has IOC problem. Beam gone.
2:39 Beam on the way.
2:42 After spending rather a lot of time trying to get replay going again, Pete had only temporary success. About 10 minutes after he went back to bed, replay quit working. I think at this point we will just do without. This means we are doing without real time data monitoring but we will keep a close eye on the various systems
4:41 During shift checklist found that dead time was high: 25%. Don't know if this is due to problems since we can't look at replay histos. All systems seem to be operating normally.
6:30 Pete called about dead time problem. We were getting error messages on ROC7 (no data in TDC). Tried rebooting, no luck. Tried different trigger (withrich), no luck. Pete thought we might have a network problem so we paged Bob Michaels. At his suggestion we set all prescale values high except the pulser (#8) and got no deadtime. Also did a copy between adaql2 and adaql1 and found fast transfer time. According to Bob, network is probably not the problem. He is on his way in for 8:00 shift. We will run as is until he gets here.
6:51 Going back to original trigger configuration (richf1) and keep running. Runs 4803-4809 are junk. Runs from 4797 up may also be bad.
Going to try a reboot...no effect--still have 29% dead time.
7:35 Pete's here looking at problem. Bob's on his way.